The altitude of an isosceles triangle is 15 centimeters. The base angles measure 35 degrees. Find the length of the base of the Triangle.

Answer by ankor@dixie-net.com(22740)

The altitude of an isosceles triangle is 15 centimeters.
The base angles measure 35 degrees.
Find the length of the base of the Triangle.
:
We can find the half length (x) of the base using the tangent of 35 degrees
tan(35) = 15/x
x = 21.42 * 2 = 42.84 is the length of the base